### v1

# 1.1 Preliminaries on probes

1 Quantum and Classical Probes over a Channel
In this section, we consider quantum and classical probes that are sent through a lossy channel with transmissivity $\eta_{0}\in(0,1)$. Suppose at an unknown time $\nu$, the transmissivity of the channel drops from $\eta_{0}=\eta\in(0,1)$ to $\eta_{1}=\eta_{0}\eta_{d}$, $\eta_{d}\in(0,1)$. We first review existing results from [guha2025quantum] in § 1.1, and then proceed to derive the quantum advantage in reducing detection delay over a channel in § 1.2.
1.1 Preliminaries on probes
We begin with a brief introduction to the quantum and classical probes used in this paper, followed by a short description of how the probes are used to detect transmissivity change in a channel.
Pre-change and post-change distributions.
As in [guha2025quantum], the classical (or quasi-classical) benchmark we consider is a stream of $T$ coherent-state pulses |$\ket{\alpha}^{\otimes T}$, with $\alpha\in\mathbb{R}$ and mean photon number $\alpha^{2}=N+N_{a}$. Here $0<N_{a}<<N$ is introduced to fairly compare the classical case with its quantum-augmented counterparts. At the receiving monitor, |$\ket{\alpha}^{\otimes T}$ generates $T$ i.i.d. Gaussian-distributed random variables $X_{1},X_{2},\dots,X_{T}$, where
$X_{t}\sim\mathcal{N}^{(j)}(\sqrt{\eta_{j}}\alpha,\,\frac{1}{4}),$ (1)
and $j=0$ for all $0\leq t<\nu$ (pre-change), $j=1$ for all $\nu\leq t\leq T$ (post-change).
Next, we consider the entanglement-augmented probes. Specifically, $T$ blocks of $n\in\mathbb{Z}_{+}$ coherent-state pulses |$\ket{\alpha}^{\otimes n}$ are augmented by a continuous variable entangled state generated by splitting a squeezed-vaccum state |$\ket{0;s}$, $s\in\mathbb{R}_{+}$, where
$N=\alpha^{2},$ (2)
$\sinh^{2}(s)=nN_{a}.$ (3)
In this case, the receiving monitor outputs $T$ i.i.d. $n$-dimensional Gaussian random vectors $\bm{X}_{1},\bm{X}_{2},\dots,\bm{X}_{T}$, where
$\bm{X}_{t}\sim\mathcal{N}_{n}^{(j)}(\bm{\mu}^{(j)},\,\bm{\Sigma}^{(j)}),$ (4)
and $j=0$ pre-change, $j=1$ post-change. Denote $\bm{I}$ as the $n\times n$ identity matrix, $\bm{J}$ as the $n\times n$ matrix of ones, and $\bm{u}$ as the $n$-dim column vector of ones. The mean vectors $\bm{\mu}^{(j)}$ and the covariance matrices $\bm{\Sigma}^{(j)}$ are given by
$\bm{\mu}^{(j)}=\sqrt{\eta_{j}}\alpha\cdot\bm{u},$
$\bm{\Sigma}^{(j)}=\frac{1}{4}\bm{I}-\frac{\eta_{j}(1-e^{-2s})}{4n}\cdot\bm{J}.$ (5)
In particular, when $n=1$, we have a stream of $T$ displaced squeezed states |$\ket{\alpha;r}^{\otimes T}$, with $\alpha^{2}=N$ and $\sinh^{2}(r)=N_{a}$. The receiving monitor outputs $T$ i.i.d. Gaussian-distributed random variables $X_{1},X_{2},\dots,X_{T}$, where
$X_{t}\sim\mathcal{N}_{1}^{(j)}(\sqrt{\eta_{j}}\alpha,\,\sigma_{j}^{2}),$ (6)
and $\sigma_{j}^{2}=\frac{1}{4}(\eta_{j}e^{-2r}+1-\eta_{j})$, and $j=0$ pre-change, $j=1$ post-change. While the physical processes of generating the probes for $n=1$ and $n\geq 2$ are different [guha2025quantum], mathematically, there is in fact no need to distinguish these two cases. We only make the distinction when discussing previous results in relation to ours in § 1.2. Whenever necessary, we refer to $n=1$ as the squeezing augmentation case, and $n\geq 2$ the general quantum augmentation case.
Quickest change detection.
Let $X_{1},X_{2},\dots,X_{T}$ be the observations from any type of probes introduced above. Suppose a transmission drop occurs at an unknown timestep $0<\nu\leq T$, and $f_{0}(x),f_{1}(x)$ are the pre-change and post-change distributions respectively (e.g., (1) for classical probes, (4) for quantum probes). To detect the change point $\nu$, the cumulative sum (CUSUM) algorithm keeps a counter $\text{CUSUM}[t]$ that can be updated recursively using
$\text{CUSUM}[t]=\max\left\{0,\text{CUSUM}[t]+\log{\frac{f_{1}(X_{t})}{f_{0}(X_{t})}}\right\},\forall t\in[T],$
and $\text{CUSUM}[0]=0$. Let $\gamma$ be the minimum time to false alarm that one is willing to tolerate, and denote $\tau$ as the minimum $t$ such that $\text{CUSUM}[t]>\log{\gamma}$. Then the CUSUM algorithm outputs $\tau$ as the estimate for the change point $\nu$. Informally, as $\gamma\to\infty$, the CUSUM algorithm achieves a worst-case detection delay [L1971] of
$\frac{\log{\gamma}}{D(f_{1}||f_{0})}(1+o(1)),$ (7)
where $D(f_{1}||f_{0})=\int_{-\infty}^{\infty}f_{1}(x)\log{\left(\frac{f_{1}(x)}{f_{0}(x)}\right)}\,dx$ is the Kullback-Leibler (KL) divergence from $f_{0}$ to $f_{1}$.
Kullback-Leibler (KL) divergence.
Given two $n$-dimensional Gaussian distributions $\mathcal{N}_{n}^{(j)}(\bm{\mu}^{(j)},\,\bm{\Sigma}^{(j)})$, $j=0,1$, with non-singular covariance matrices $\bm{\Sigma}^{(j)}$ and $\bm{\Sigma}^{(j)}$, the KL divergence from $\mathcal{N}_{n}^{(0)}$ to $\mathcal{N}_{n}^{(1)}$ is
$D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)=\frac{1}{2}\left(\operatorname{Tr}\left({\left(\bm{\Sigma}^{(0)}\right)^{-1}\bm{\Sigma}^{(1)}}\right)+\left(\bm{\mu}^{(0)}-\bm{\mu}^{(1)}\right)^{T}\left(\bm{\Sigma}^{(0)}\right)^{-1}\left(\bm{\mu}^{(0)}-\bm{\mu}^{(1)}\right)-n+\ln{\frac{\left\lvert{\bm{\Sigma}^{(0)}}\right\rvert}{\left\lvert{\bm{\Sigma}^{(1)}}\right\rvert}}\right),$ (8)
where $\left\lvert{\bm{\Sigma}^{(j)}}\right\rvert$ denotes the determinant of $\bm{\Sigma}_{j}$, $j=0,1$. By (1) and (8) in the case of $n=1$, the KL divergence for each coherent state pulse is
$D(\mathcal{N}^{(1)}||\mathcal{N}^{(0)})=2(N+N_{a})(\sqrt{\eta_{0}}-\sqrt{\eta_{1}})^{2}=2(N+N_{a})\eta(1-\sqrt{\eta_{d}})^{2}.$ (9)
To get the per-pulse KL divergence for entanglement-augmented coherent states, we take the average over $n$ pulses in that block. That is, we use $\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)$ as a fair comparison to the classical case (9), with $D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)$ as shown in (8).
1.2 The quantum advantage in reducing detection delay
Since the worst-case detection delay is inversely proportional to the KL divergence (7), to quantify how quantum probes fare against the classical benchmark in speeding up the change point detection, we consider the ratio of quantum to classical KL divergences. To start with, we derive the explicit form of this ratio as a function of $\eta\in(0,1)$, $\eta_{d}\in(0,1)$, $N\in\mathbb{Z}_{+}$ and $N_{a}\in\mathbb{Z}_{+}$, parametrized by $n\in\mathbb{Z}_{+}$.
Lemma 1.
Let $q_{n}=q_{n}(\eta,\eta_{d},N,N_{a})=\frac{D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)}{n\cdot D(\mathcal{N}^{(1)}||\mathcal{N}^{(0)})}$ be the ratio of quantum to classical KL divergences, and denote $c_{n}=\frac{2\sqrt{nN_{a}}}{\sqrt{nN_{a}+1}+\sqrt{nN_{a}}}$,
$q_{n}=\frac{1}{4(N+N_{a})n}\left(\frac{c_{n}}{1-c_{n}\eta}\cdot\frac{1-\eta_{d}}{(1-\sqrt{\eta_{d}})^{2}}-\frac{1}{\eta}\cdot\frac{\ln{\frac{1-c_{n}\eta\eta_{d}}{1-c_{n}\eta}}}{(1-\sqrt{\eta_{d}})^{2}}+\frac{4Nn}{1-c_{n}\eta}\right).$ (10)
Proof.
First note that from (5), covariance matrix $\bm{\Sigma}^{(j)}$ has uniform diagonal entries $d_{j}=\frac{1}{4}\left(1-\frac{\eta_{j}(1-e^{-2s})}{4n}\right)$ and uniform off-diagonal entries $r_{j}=d_{j}-\frac{1}{4}$. Such a structure allows us to explicitly write out both $\left\lvert{\bm{\Sigma}^{(j)}}\right\rvert$ and $\left(\bm{\Sigma}^{(j)}\right)^{-1}$.
$\left\lvert{\bm{\Sigma}^{(j)}}\right\rvert=\begin{vmatrix}d_{j}+(n-1)r_{j}&d_{j}+(n-1)r_{j}&\cdots&d_{j}+(n-1)r_{j}\\ r_{j}&d_{j}&\cdots&r_{j}\\ \vdots&\vdots&\ddots&\vdots\\ r_{j}&r_{j}&\cdots&d_{j}\end{vmatrix}=(\frac{1}{4}+nr_{j})\cdot\frac{1}{4^{n-1}}.$
Hence the last term in (8) is
$\ln{\frac{\left\lvert{\bm{\Sigma}^{(0)}}\right\rvert}{\left\lvert{\bm{\Sigma}^{(1)}}\right\rvert}}=-\ln{\frac{1-\eta\eta_{d}(1-e^{-2s})}{1-\eta(1-e^{-2s})}}.$ (11)
By (3), $e^{-2s}=1-2\sqrt{nN_{a}}(\sqrt{nN_{a}+1}-\sqrt{nN_{a}})$, and
$1-e^{-2s}=\frac{2\sqrt{nN_{a}}}{\sqrt{nN_{a}+1}+\sqrt{nN_{a}}}\triangleq c_{n}.$ (12)
Substitution of (12) into (11) yields
$\ln{\frac{\left\lvert{\bm{\Sigma}^{(0)}}\right\rvert}{\left\lvert{\bm{\Sigma}^{(1)}}\right\rvert}}=-\ln{\frac{1-c_{n}\eta\eta_{d}}{1-c_{n}\eta}}.$ (13)
To derive $\left(\bm{\Sigma}^{(0)}\right)^{-1}$, observe that $\bm{\Sigma}^{(j)}$ (5) is the sum of a diagonal matrix and a rank-$1$ matrix, that is, $\bm{\Sigma}^{(j)}=\frac{1}{4}\bm{I}+\bm{u}(r_{j}\bm{u})^{T}$. By the Sherman-Morrison formula,
$\left(\bm{\Sigma}^{(0)}\right)^{-1}=\left(\frac{1}{4}\bm{I}\right)^{-1}-\frac{\left(\frac{1}{4}\bm{I}\right)^{-1}\cdot\bm{u}(r_{0}\bm{u})^{T}\cdot\left(\frac{1}{4}\bm{I}\right)^{-1}}{1+(r_{0}\bm{u})^{T}\left(\frac{1}{4}\bm{I}\right)^{-1}\bm{u}}=4\bm{I}+\frac{4c_{n}\eta}{n(1-c_{n}\eta)}\bm{J}.$
Then we are ready to write out the first two term in (8),
$\operatorname{Tr}\left({\left(\bm{\Sigma}^{(0)}\right)^{-1}\bm{\Sigma}^{(1)}}\right)=n+\frac{c_{n}\eta}{1-c_{n}\eta}(1-\eta_{d}),$ (14)
$\left(\bm{\mu}^{(0)}-\bm{\mu}^{(1)}\right)^{T}\left(\bm{\Sigma}^{(0)}\right)^{-1}\left(\bm{\mu}^{(0)}-\bm{\mu}^{(1)}\right)=(\sqrt{\eta_{0}}-\sqrt{\eta_{1}})^{2}\cdot\alpha^{2}\cdot\bm{u}^{T}\left(\bm{\Sigma}^{(0)}\right)^{-1}\bm{u}=\frac{4Nn\eta(1-\sqrt{\eta_{d}})^{2}}{1-c_{n}\eta},$ (15)
where the last equality in (15) uses (2). Substituting (14), (15) and (13) into (8), we arrive at the final form of $\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)$,
$\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)=\frac{1}{2n}\left(\frac{c_{n}\eta}{1-c_{n}\eta}(1-\eta_{d})-\ln{\frac{1-c_{n}\eta\eta_{d}}{1-c_{n}\eta}}+\frac{4Nn\eta}{1-c_{n}\eta}(1-\sqrt{\eta_{d}})^{2}\right).$ (16)
Therefore, the form of $q_{n}$ in (10) follows from dividing the per-pulse quantum KL divergence (16) by the classical KL divergence (9). ∎
Using the explicit form of the per-pulse quantum KL divergence (16), we are able to rigorously derive several results that have been previously inferred through numerical evaluations in [guha2025quantum]. It has been shown that the KL divergence sees a sharp increase when augmenting a tiny amount of squeezing. That is, when $n=1$, $\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)=\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}D\left(\mathcal{N}_{1}^{(1)}||\mathcal{N}_{1}^{(0)}\right)=\infty$. It has also been observed without proof that the same holds for general $n$. By taking the partial derivative of (16), it is straightforward to verify that indeed,
$\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)=2N\sqrt{n}\eta^{2}(1-\sqrt{\eta_{d}})^{2}\lim_{N_{a}\to 0}\frac{1}{\sqrt{N_{a}}}=\infty.$
Additionally, numerical evaluations (e.g. Figure 2 in [guha2025quantum], with $n=2^{i}$, $i\in[8]$) suggest that for fixed $N$, $N_{a}$, $\eta$ and $\eta_{d}$, the quantum KL divergence $\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)$ converges to $\frac{2N\eta(1-\sqrt{\eta_{d}})^{2}}{1-\eta}$ as $n$ becomes large. This directly follows from taking the limit of (16) as $n$ goes to infinity.
With Lemma 1 in hand, we are well-equipped to understand, given a set of parameters, whether quantum probes outperform the classical ones, and if so, by how much. Moreover, we can explore how the quantum advantage changes with respect to each variable, and what is the asymptotic behavior of such advantage.
Lemma 2.
$q_{n}=q_{n}(\eta,\eta_{d},N,N_{a})$ exhibits the following properties:
  1. (1)
    For fixed values of $n$, $\eta_{d}$, $N$ and $N_{a}$, $q_{n}$ increases in $\eta$, and
    $\lim_{\eta\to 0}q_{n}=\frac{N}{N+N_{a}};$ (17)
  2. (2)
    For fixed values of $n$, $\eta$, $N$ and $N_{a}$, $q_{n}$ increases in $\eta_{d}$, and
    $\lim_{\eta_{d}\to 1}q_{n}=\frac{1}{2n(N+N_{a})(1-c_{n}\eta)}\left(\frac{c_{n}^{2}\eta}{1-c_{n}\eta}+2nN\right);$ (18)
  3. (3)
    For fixed values of $n$, $\eta$, $\eta_{d}$ and $N_{a}$, $q_{n}$ increases in $N$ when $N>\frac{c_{n}^{2}t(1+\sqrt{\eta_{d}})^{2}}{4nN_{a}(1-c_{n}\eta\eta_{d})}$, and
    $\lim_{N\to\infty}q_{n}=\frac{1}{1-c_{n}\eta};$ (19)
  4. (4)
    For fixed values of $\eta$, $\eta_{d}$, $N$ and $N_{a}$ satisfying $N\eta>b_{d}N_{a}\left(1-\eta\right)$ with $b_{d}=\frac{1+\sqrt{\eta_{d}}}{1-\sqrt{\eta_{d}}}$, $q_{n}$ increases in $n$ when $n\geq n_{0}$, where
    $n_{0}=\begin{cases}\max\left(\frac{b_{d}\left(3-4\eta\right)}{4\left(N\eta-b_{d}N_{a}\left(1-\eta\right)\right)},\frac{1}{4N_{a}(1-\eta)}\right)&\text{if }0<\eta<\frac{1}{2},\\ \max\left(\frac{b_{d}}{4\left(N\eta-b_{d}N_{a}\left(1-\eta\right)\right)},\frac{1}{4N_{a}(1-\eta)}\right)&\text{if }\frac{1}{2}\leq\eta<1,\end{cases}$ (20)
    and
    $\lim_{n\to\infty}q_{n}=\frac{N}{(N+N_{a})(1-\eta)};$ (21)
  5. (5)
    For fixed values of $n$, $\eta$, $\eta_{d}$ and $N$:
    1. (5.1)
      For any $\varepsilon>0$, denote $c_{n,\varepsilon}=\frac{2\sqrt{n\varepsilon}}{\sqrt{n\varepsilon+1}+\sqrt{n\varepsilon}}$, if $\varepsilon\leq N_{a}\leq\frac{c_{n,\varepsilon}N\eta}{1-c_{n,\varepsilon}\eta}$, we have $q_{n}>1$;
    2. (5.2)
      $\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}q_{n}=\sqrt{n}\eta\lim_{N_{a}\to 0}\frac{1}{\sqrt{N_{a}}}=\infty;$
    3. (5.3)
      If $8Nn\left(1-\eta\right)>\eta b_{d}\left(1-\eta_{d}\right)$ with $b_{d}$ defined as above, $q_{n}$ decreases in $N_{a}$ when $N_{a}\geq N_{a,0}$, where
[…]
Lemma 2(1)(3) not only corroborates an observation that quantum augmentation provides the most benefit in the large-$N$ and high-$\eta$ regime [guha2025quantum] when $n=1$, but also shows that the same holds in the general quantum augmentation case (arbitrary $n$). Additionally, [guha2025quantum] provides an approximate threshold $\frac{N\eta}{1-\eta}$ on $N_{a}$, above which squeezing augmentation does not outperform the classical benchmark in the case of $n=1$. Lemma 2(5)(5.1) gives a range of $N_{a}$ where quantum advantage exists, for any $n$. We note that, for typical values of $N$, $n$, $\eta$ and $\eta_{d}$ (e.g., see § LABEL:sec:simulation and [guha2025quantum]), $q_{n}$ is likely to first increase in $N_{a}$, and then decrease. Lacking ways to directly find the $N_{a}$ that maximizes $q_{n}$, we settle for the result in Lemma 2(5)(5.3). Nonetheless, $N_{a,0}$ may be of practical interest, as it narrows down the range of $N_{a}$ in which the optimal quantum advantage appears. Too see that, given typical values such as $N=100$, $n=1$, $\eta=0.9$ and $\eta_{d}=0.8$, Lemma 2(5)(5.1) gives that the quantum advantage exists when $897.50<N_{a}<$ 897.5 by taking $\epsilon$ ranging from arbitrarily close to 00 to $890$, and [guha2025quantum] shows that such advantage disappears when $N_{a}$ is above roughly $900$; however, both bounds on $N_{a}$ are rendered irrelevant under the assumption that $N_{a}<<N=100$. In contrast, Lemma 2(5)(5.3) indicates that quantum augmentation brings the most value when $N_{a}<N_{a,0}=20.9$. This is evidence that assuming $N_{a}<<N$ (§ 1.1) makes sense.
Beyond generalizing and strengthening previous results, we can draw a few interesting conclusions from Lemma 2. While smaller transmissivity drops (larger $\eta_{d}$) are generally harder to detect, the quantum advantage increases as the drop gets lighter (Lemma 2(2)). Moreover, with proper choices of parameters, the quantum advantage exists, and the limits in Lemma 2(2)(3)(4) provide the respective upper bounds on how much quantum augmentation could help.
In spite of the generality of Lemma 2, we acknowledge that the lower bounds of $n$ (in Lemma 2(4)) and $N_{a}$ (in Lemma 2(5)(5.1)) are likely consequences of our proof technique, rather than being necessary. We leave further improvements for future work.